# GraphGPT: Generative Pre-trained Graph Eulerian Transformer

### v2

@@ Abstract
We introduce GraphGPT, a novel self-supervised generative pre-trained model for graph learning based on the Graph Eulerian Transformer (GET). First, we propose GET, which combines a standard transformer encoder or decoder architecture with an innovative graph-to-sequence transformation method. This method converts graphs or sampled subgraphs into sequences of tokens representing nodes, edges, and attributes in a reversible manner using Eulerian paths. We pre-train GET using either of the two self-supervised tasks: next-token prediction (NTP) and scheduled masked-token prediction (SMTP). The pre-trained model is then fine-tuned for downstream tasks such as graph-, edge-, and node-level prediction. Despite its simplicity, GraphGPT achieves performance comparable to or surpassing state-of-the-art methods on multiple large-scale Open Graph Benchmark (OGB) datasets. It demonstrates exceptional results on the molecular property prediction dataset PCQM4Mv2 and the protein-protein interaction dataset ogbl-ppa. Notably, generative pre-training enables scaling GraphGPT to 2 billion parameters while maintaining performance gains — a breakthrough that overcomes the scalability limitations of traditional Graph Neural Networks (GNNs) and prior graph transformers (GTs). To advance research in graph foundation models and facilitate scientific discovery in chemistry, materials science, and related fields, we will release the source code and pre-trained checkpoints.
@@ 1 Introduction
Despite progress, GNNs remain constrained by over-smoothing (Rusch et al., 2023) and over-squashing (Alon & Yahav, 2021), limiting their scalability and capacity to leverage large-scale graph data. Recent efforts to adapt transformers to graphs (Ying et al., 2021; Kim et al., 2022; Luo et al., 2023; Müller et al., 2023) show promise but face critical challenges: 1). Structural Bias: Most graph transformers (GTs) rely on handcrafted features or GNN modules to encode graph topology, compromising generalization. 2). Task Limitations: GTs excel at graph-level tasks but struggle with edge- and node-level objectives (Müller et al., 2023). 3). Pre-Training Gap: Unlike NLP’s success with self-supervised pre-training (Radford et al., 2018; Devlin et al., 2019), GTs lack effective frameworks for generative pre-training (Min et al., 2022; Müller et al., 2023).
@@
Graph Eulerian Transformer (GET): We introduce GET, a novel architecture that leverages (semi-)Eulerian paths111For a quick recap, a graph with every node of even degree is Eulerian, and with exactly two nodes of odd degree is semi-Eulerian. to losslessly and reversibly convert graphs into token sequences. By integrating subgraph sampling and node identity encoding, GET efficiently processes graphs of arbitrary sizes. A standard transformer encoder or decoder is then applied to these sequences, eliminating the need for specialized architectural modifications.
@@ 2.2 Graph to Sequence of Tokens
Large graphs (with up to billions of nodes/edges and rich attributes) are first decomposed into subgraphs via the sampling process described in §2.2.2. Node identity preservation (§2.2.3) ensures structural consistency during this decomposition. These subgraphs are then serialized using §2.2.1’s guidelines.
@@ 2.2.1 Serializing graphs with (semi-)Eulerian Paths
2.
Bijective mapping between the graph and its serialized form.
@@
Implementation Steps:
1.
Path Identification:
@@
Randomly sample one valid path from possible candidates, introducing stochasticity as a data augmentation strategy akin to computer vision techniques (Perez & Wang, 2017).
2.
Node Re-indexing:
@@
Introduce cyclic re-indexing: , where is a random integer and (hyperparameter) exceeds the maximum node count. This ensures uniform token training by distributing index frequencies.
3.
Attribute Handling:
@@
4.
Disconnected Components:
Connect components by adding synthetic edges between randomly selected nodes.
Theoretical Guarantee: The serialization is lossless and reversible (up to isomorphism) per the Eulerian Path Theorem: reconstructing edges from adjacent tokens recovers the original graph structure (Grohe & Schweitzer, 2020). For example, in Fig. 1’s sequence, connecting consecutive tokens () reconstructs all edges, yielding a graph isomorphic to the input.
@@ 2.2.2 Subgraph Sampling
1.
Computational Overhead: Eulerization and path identification for massive graphs are computationally prohibitive.
2.
@@
To address these challenges, we adopt subgraph sampling—a scalable strategy that decomposes large graphs into smaller, manageable subgraphs before serialization.
@@ 2.2.3 Node Identity Encoding
Our ablation studies (§3.5.2) demonstrate this method’s effectiveness in preserving node identity while maintaining tractable model scaling.
@@ 2.3.1 Pre-training with the NTP or SMTP tasks
Implementation Details:
Masking (SMTP): For node-level masking, all occurrences of a masked node in the Eulerian sequence are hidden to prevent leakage (e.g., two occurrences of node in Fig. 1 are masked concurrently).
Mask Scheduling (SMTP): Following Chang et al. (2022), we use a linear scheduling function, which empirically balances training stability and performance.
@@ 2.3.2 Fine-tuning on downstream graph tasks
The final token’s output is fed to a randomly initialized multilayer perceptron (MLP).
Parameter Update:
Transformer weights are initialized from pre-trained checkpoints.
MLP layers are trained from scratch.
All parameters are updated during fine-tuning.
@@ 3.2 Graph-level tasks
| Models | MAE | Params | |
| Valid | Test | ||
| GCN | 0.1379 | 0.1398 | 2.0M |
| GIN | 0.1195 | 0.1218 | 3.8M |
| GCN-VN | 0.1153 | 0.1152 | 4.9M |
| GIN-VN | 0.1083 | 0.1084 | 6.7M |
| TokenGT | 0.0910 | 0.0919 | 48.5M |
| Graphformer | 0.0864 | N/A | 48.3M |
| GPS-Deep | 0.0852 | 0.0862 | 138.1M |
| GPS++ (no 3D) | 0.0818 | N/A | 40.0M |
| GPTrans-L | 0.0809 | 0.0821 | 86.0M |
| GraphGPT-B12 | 0.0807 | N/A | 113.6M |
| GraphGPT-B24 | 0.0793 | N/A | 227.3M |
| GraphGPT-B48 | 0.0792 | 0.0804 | 453.4M |
We evaluate GraphGPT on two molecular datasets where tasks involve predicting quantum chemical properties solely from 2D molecular graphs—a practical alternative to relying on 3D equilibrium structures. Specifically: PCQM4Mv2 predicts the HOMO-LUMO energy gap, and ogbg-molpcba predicts 128 binary molecular properties.
@@
Scalability: While GTs with fewer parameters often plateau when scaled (Shi et al., 2022), GraphGPT shows consistent improvement up to 400M parameters.
Parameter Efficiency: GraphGPT’s larger parameter count may reflect its capacity to implicitly learn features that other GTs encode manually. Generative pre-training also allocates model capacity to generation tasks, potentially limiting discriminative performance at smaller scales.
Limitations: Pre-training with additional SMILES-formatted molecular data yielded diminishing returns, suggesting saturation in 2D structural information. Incorporating 3D molecular data could address this.
@@ 3.4 Node-level tasks
ogbn-proteins: GraphGPT surpasses well-tuned GNN baselines (GCN, GraphSAGE, GAT) and significantly outperforms graph transformers (GTs). Remarkably, GraphGPT achieves competitive performance with input subgraphs of nodes, while SOTA GNNs like AGDN (Sun et al., 2020) require subgraphs with nodes.
@@
The strong performance with minimal neighborhood sampling suggests that generative pre-training effectively encodes global structural and semantic graph information into node token embeddings and transformer parameters. This contrasts with traditional GNNs, which rely on extensive local aggregation for feature propagation.
@@ 3.5 Ablation Study
We analyze the impact of two core components of GraphGPT: pre-training and node identity encoding.
@@ 3.5.2 Node Identity Encoding
Node identity encoding (see §2.2.3)—representing nodes in large graphs as unique multi-token sequences—is critical for edge- and node-level tasks. Using GraphGPT-mini (a lightweight variant to conserve computational resources), we demonstrate that this method significantly enhances performance (Table 7). Further implementation details are provided in Appendices F and A.
@@ 4 Limitations
Computational Cost. Pre-training on large-scale graphs (ogbn-proteins, ogbl-ppa) or extensive small graphs (PCQM4Mv2) with 50M+ parameters is resource-intensive. For example, pre-training GraphGPT-B (100M+ parameters) on PCQM4Mv2 with tokens requires V100 GPU hours, and fine-tuning incurs V100 GPU hours per epoch.
While GraphGPT is less practical for small datasets due to compute-performance trade-offs, it excels with large-scale data. Emerging techniques like quantization (Dettmers et al., 2022; Frantar et al., 2022), distributed training frameworks (Rasley et al., 2020; Shoeybi et al., 2019), and transformer optimizations (Dao, 2023) are poised to alleviate these costs.
@@ 5 Related works
Graph Transformers (GTs) Inspired by transformers’ success in NLP and CV, recent work has adapted these architectures to graphs (Ying et al., 2021; Rampásek et al., 2022; Müller et al., 2023). While GTs achieve competitive results on large-scale graph-level tasks (Müller et al., 2023), they typically rely on handcrafted structural features or GNN-based modules to encode graph topology—either in input representations (Ying et al., 2021; Kim et al., 2022; Masters et al., 2023) or attention mechanisms (Ying et al., 2021; Chen et al., 2022; Luo et al., 2023).
Pre-training and fine-tuning The self-supervised pre-training and supervised fine-tuning paradigm, popularized by transformers (Vaswani et al., 2017), revolutionized NLP (Devlin et al., 2019; Radford et al., 2018). Scaling this approach with web-scale data (Brown et al., 2020) and techniques like instruction tuning (Wei et al., 2021) or reinforcement learning from human feedback (Ouyang et al., 2022) further advanced the field. In CV, self-supervised methods like MAE He et al. (2022) and MaskGIT (Chang et al., 2022)demonstrated that masked prediction tasks (e.g., reconstructing masked image patches) enable transformers to achieve SOTA results.
@@ C.1 Graphs to Sequences of tokens
The code is implemented with Pytorch. We use torch-geometric (Fey & Lenssen, 2019) to preprocess graphs, e.g., subgraph sampling and etc. We use Networkx (Hagberg et al., 2008) to Eulerize the (sub)graphs if needed and then find the (semi-)Eulerian paths. We write our own tokenizer to convert the (semi-)Eulerian paths to sequences of tokens. The token vocabulary is built for each dataset separately.
@@ C.2 Model backbone
We adopt Llama’s transformer architecture (Touvron et al., 2023) implemented in the Hugging Face’s transformers package (Wolf et al., 2020) as our backbone for the NTP pre-training. To turn it into an encoder for the SMTP pre-training, we change the causal attention mask to bi-directional attention mask. We do not use Llama’s pre-trained weights. Instead, we train models of different scales as in Tab. 10 with random parameters initialization.
@@ C.3 Training
The models are pre-trained and fine-tuned on GPU clusters of A800-80G using DeepSpeed’s stage-2 schema with mixed precision (Rasley et al., 2020) or bf16. We use AdamW (Loshchilov & Hutter, 2017) optimizer with learning rate scheduler. To make full use of computing power, we pack several graph sequences together into one entry to maximize the context window (Raffel et al., 2020). The detailed configurations for each dataset are set separately and can be found in the corresponding sections in the following.

### v3

@@ Abstract
We introduce GraphGPT, a novel self-supervised generative pre-trained model for graph learning based on the Graph Eulerian Transformer (GET). First, we propose GET, which combines a standard transformer encoder or decoder architecture with an innovative graph-to-sequence transformation method. This method converts graphs or sampled subgraphs into sequences of tokens representing nodes, edges, and attributes in a reversible manner using Eulerian paths. We pre-train GET using either of the two self-supervised tasks: next-token prediction (NTP) and scheduled masked-token prediction (SMTP). The pre-trained model is then fine-tuned for downstream tasks such as graph-, edge-, and node-level prediction. Despite its simplicity, GraphGPT achieves performance comparable to or surpassing state-of-the-art methods on multiple large-scale Open Graph Benchmark (OGB) datasets. It demonstrates exceptional results on the molecular property prediction dataset PCQM4Mv2 and the protein-protein interaction dataset ogbl-ppa. Notably, generative pre-training enables scaling GraphGPT to 2 billion parameters while maintaining performance gains — a breakthrough that overcomes the scalability limitations of traditional Graph Neural Networks (GNNs) and prior graph transformers (GTs). To advance research in graph foundation models and facilitate scientific discovery in chemistry, materials science, and related fields, we have released the source code111https://github.com/alibaba/graph-gpt and model checkpoints222https://www.modelscope.cn/organization/Alibaba-DT.
@@ 1 Introduction
Despite progress, GNNs remain constrained by over-smoothing (Rusch et al., 2023) and over-squashing (Alon & Yahav, 2021), limiting their scalability and capacity to leverage large-scale graph data. Recent efforts to adapt transformers to graphs (Ying et al., 2021; Kim et al., 2022; Luo et al., 2023; Müller et al., 2024) show promise but face critical challenges: 1). Structural Bias: Most graph transformers (GTs) rely on handcrafted features or GNN modules to encode graph topology, compromising generalization. 2). Task Limitations: GTs excel at graph-level tasks but struggle with edge- and node-level objectives (Müller et al., 2024). 3). Pre-Training Gap: Unlike NLP’s success with self-supervised pre-training (Radford et al., 2018; Devlin et al., 2019), GTs lack effective frameworks for generative pre-training (Min et al., 2022; Müller et al., 2024).
@@
Graph Eulerian Transformer (GET): We introduce GET, a novel architecture that leverages Eulerian or semi-Eulerian paths333For a quick recap, a connected graph with every node of even degree is Eulerian, and with exactly two nodes of odd degree is semi-Eulerian. In this paper, ‘(semi-)Eulerian’ refers to both Eulerian and semi-Eulerian unless specified otherwise. to losslessly and reversibly convert graphs into token sequences. By integrating subgraph sampling and node identity encoding, GET efficiently processes graphs of arbitrary sizes. A standard transformer encoder or decoder is then applied to these sequences, eliminating the need for specialized architectural modifications.
@@ 2.2.1 Serializing Graphs with (Semi-)Eulerian Paths
Bijective mapping between the graph and its serialized form444Given a fixed starting node and predetermined node indices, the Eulerian path generated by NetworkX (Hagberg et al., 2008) is guaranteed to be unique..
@@
Implementation:
1. Connectivity Check:
Check if the graph is connected. If not, link disconnected components by adding synthetic edges between randomly selected nodes. For example, given a graph with disconnected components A, B, and C, connect A and B via a random node pair, then B and C similarly.
This ensures the graph becomes connected, enabling Eulerian path generation.
2. Path Identification:
@@
Randomly sample one valid path from possible candidates, introducing stochasticity as a data augmentation strategy akin to computer vision techniques (Perez & Wang, 2017). This stochasticity forces the transformer to learn invariance across different paths of the same graph and empirically reduces overfitting.
3. Node Re-indexing:
@@
Introduce cyclic re-indexing: , where is a random integer and (hyperparameter) exceeds the maximum node count. Without cyclic re-indexing, Eulerian paths would always start with low-index tokens (e.g., ), leading to skewed token frequency distributions. Cyclic re-indexing randomizes starting indices (e.g., selecting from for ), ensuring uniform training across all index tokens.
Cyclic re-indexing is critical for datasets like Triangles (§3.2.1), where test graphs have significantly more nodes than training graphs (e.g., test graphs up to nodes vs. training graphs nodes). Without re-indexing, higher-index tokens (e.g., ) remained undertrained, degrading performance.
4. Attribute Attaching:
@@
Theoretical Guarantee:
The serialization is lossless and reversible (up to isomorphism) per the Eulerian Path Theorem: reconstructing edges from adjacent tokens recovers the original graph structure (Grohe & Schweitzer, 2020). For example, in Fig. 1’s sequence, connecting consecutive tokens () reconstructs all edges, yielding a graph isomorphic to the input.
@@ 2.2.2 Subgraph Sampling
Computational Overhead: Eulerization and path identification for a large graph are computationally expensive.
@@
To address these challenges, we adopt subgraph sampling—a scalable strategy that decomposes a large graph into smaller, manageable subgraphs before serialization.
@@ 2.2.3 Node Identity Encoding
Our ablation studies (§3.5.3) demonstrate this method’s effectiveness in preserving node identity.
@@ 2.3.1 Pre-training with the NTP or SMTP Tasks
Implementation:
Masking (SMTP only): For node-level masking, all occurrences of a masked node in the Eulerian sequence are hidden to prevent leakage (e.g., two occurrences of node in Fig. 1 are masked concurrently).
Mask Scheduling (SMTP only): Following Chang et al. (2022), we adopt the same linear scheduling function, which empirically balances training stability and performance.
@@ 2.3.2 Fine-tuning on Downstream Graph Tasks
The final token’s output is fed to a randomly initialized multilayer perceptron (MLP). Fig. 2 (Appendix) illustrates the implementation.
The transformer weights are initialized from pre-trained checkpoints, and the MLP layers are initialized randomly. All parameters are updated during fine-tuning.
@@ 3.2 Graph-Level Tasks
| Models | MAE | Params | ||
| Valid | Test | |||
| GNN | GCN | 0.1379 | 0.1398 | 2.0M |
| GIN | 0.1195 | 0.1218 | 3.8M | |
| GCN-VN | 0.1153 | 0.1152 | 4.9M | |
| GIN-VN | 0.1083 | 0.1084 | 6.7M | |
| GT | TokenGT | 0.0910 | 0.0919 | 48.5M |
| Graphformer | 0.0864 | N/A | 48.3M | |
| GPS-Deep | 0.0852 | 0.0862 | 138.1M | |
| GPS++ (no 3D) | 0.0818 | N/A | 40.0M | |
| GPTrans-L | 0.0809 | 0.0821 | 86.0M | |
| Ours | GraphGPT-M | 0.0827 | N/A | 37.7M |
| GraphGPT-B12 | 0.0807 | N/A | 113.6M | |
| GraphGPT-B24 | 0.0793 | N/A | 227.3M | |
| GraphGPT-B48 | 0.0792 | 0.0804 | 453.4M | |
We evaluate GraphGPT on two molecular datasets where tasks involve predicting quantum chemical properties solely from 2D molecular graphs—a practical alternative to relying on 3D equilibrium structures. Specifically, PCQM4Mv2 predicts the HOMO-LUMO energy gap, and ogbg-molpcba predicts 128 binary molecular properties.
@@
Scalability: While GTs with fewer parameters often plateau when scaled (Shi et al., 2022), GraphGPT shows consistent improvement up to 200M parameters. The log-log scaling law plot for both pre-training loss and supervised fine-tuning loss is shown in Fig. 3 (Appendix).
Parameter Efficiency: GraphGPT’s larger parameter count may reflect its capacity to implicitly learn features that other GTs encode manually. Generative pre-training also allocates model capacity to generation tasks, potentially limiting discriminative performance of models at smaller scales.
Limitations: Pre-training on additional external large-scale molecular datasets yielded diminishing returns, suggesting saturation in 2D structural information. Incorporating 3D molecular data could help address this limitation.
@@ 3.4 Node-Level Tasks
ogbn-proteins: GraphGPT surpasses well-tuned GNN baselines (GCN, GraphSAGE, GAT) and significantly outperforms graph transformers (GTs). Remarkably, GraphGPT achieves competitive performance with input subgraphs of nodes, while SOTA GNNs like AGDN (Sun et al., 2025) require subgraphs with nodes.
@@
The strong performance with minimal neighborhood sampling suggests that generative pre-training effectively encodes global structural and semantic graph information into node token embeddings and transformer parameters. This contrasts with traditional GNNs, which rely on extensive local aggregation with feature propagation.
@@ 3.5 Ablation Study
We analyze the impact of three core components of GraphGPT: pre-training, node re-indexing and node identity encoding.
@@ 3.5.1 Pre-training
3.5.2 Node Re-indexing
As illustrated in Fig. 1, we re-index the nodes based on their order in the (semi-)Eulerian path. To evaluate the effectiveness of this approach, we conduct experiments on the ogbg-molpcba dataset, with results summarized in Tab. 7.
While node re-indexing increases pre-training loss, it consistently improves performance on downstream tasks across various model sizes. This technique acts as a form of data augmentation, preventing the model from memorizing graph-specific artifacts—such as arbitrary node labeling—and thereby enhancing generalization.
Furthermore, re-indexing enables constrained decoding of node tokens during graph generation with GraphGPT, reducing the search space for valid outputs.
| Params | Re-index | PT Loss | Test | Valid |
| 4.48M | ✗ | 0.0844 | 23.10 | 25.25 |
| ✓ | 0.0874 | 23.85 | 27.77 | |
| 114.12M | ✗ | 0.0689 | 22.70 | 26.21 |
| ✓ | 0.0750 | 25.17 | 28.57 |
@@ 3.5.3 Node Identity Encoding
Node identity encoding (see §2.2.3)—representing nodes’ identity in large graphs as multiple tokens—is critical for edge- and node-level tasks. Using GraphGPT-mini (a lightweight variant to conserve computational resources), we demonstrate that this method significantly enhances performance (Table 8). Further implementation details are provided in Appendices F and A.
@@ 4 Limitations
Computational Cost. Pre-training on large-scale graphs (ogbn-proteins, ogbl-ppa) or extensive small graphs (PCQM4Mv2) with 50M+ parameters is resource-intensive. For example, pre-training GraphGPT-B (100M+ parameters) on PCQM4Mv2 with tokens requires V100 GPU hours, and fine-tuning incurs V100 GPU hours per epoch in the distributed data parallel setting with 4 GPUs.
While GraphGPT is less practical for small datasets due to compute-performance trade-offs, it excels with large-scale data. Emerging techniques like quantization (Dettmers et al., 2022; Frantar et al., 2022), distributed training frameworks (Rasley et al., 2020; Shoeybi et al., 2019), and transformer optimizations (Dao, 2024) are poised to alleviate these costs.
@@ 5 Related Works
Graph Transformers (GTs) Inspired by transformers’ success in NLP and CV, recent work has adapted these architectures to graphs (Ying et al., 2021; Rampásek et al., 2022; Müller et al., 2024). While GTs achieve competitive results on large-scale graph-level tasks (Müller et al., 2024), they typically rely on handcrafted structural features or GNN-based modules to encode graph topology—either in input representations (Ying et al., 2021; Kim et al., 2022; Masters et al., 2023) or attention mechanisms (Ying et al., 2021; Chen et al., 2022; Luo et al., 2023).
Pre-training and fine-tuning The self-supervised pre-training and supervised fine-tuning paradigm, popularized by transformers (Vaswani et al., 2017), revolutionized NLP (Devlin et al., 2019; Radford et al., 2018). Scaling this approach with web-scale data (Brown et al., 2020) and techniques like instruction tuning (Wei et al., 2022) or reinforcement learning from human feedback (Ouyang et al., 2022) further advanced the field. In CV, self-supervised methods like MAE He et al. (2022) and MaskGIT (Chang et al., 2022) demonstrated that masked prediction tasks (e.g., reconstructing masked image patches) enable transformers to achieve SOTA results.
@@ 6 Conclusions
Acknowledgements
We thank our colleagues Guoshuai Wang and Tiange Xu for their insightful discussions and feedback. We appreciate the constructive feedback from the anonymous reviewers, which significantly improved the clarity of this paper. We thank our families for their unwavering support and encouragement throughout this project.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ C.1 Graphs to Sequences of Tokens
The implementation uses PyTorch as the primary framework. For graph preprocessing tasks such as subgraph sampling, we utilize torch-geometric (Fey & Lenssen, 2019). When required, we employ NetworkX (Hagberg et al., 2008) to Eulerize (sub)graphs and identify (semi-)Eulerian paths. A custom tokenizer converts these paths into token sequences, with dataset-specific vocabularies constructed for each case.
@@ C.2 Model Backbone
We employ a transformer architecture based on Llama (Touvron et al., 2023), implemented via the Hugging Face Transformers library (Wolf et al., 2020), as the backbone for NTP pre-training. For SMTP pre-training, we modify the architecture by replacing the causal attention mask with a bidirectional attention mask to create an encoder. We initialize all parameters randomly, and train models at various scales (see Table 11).
@@ C.3 Training
The models are pre-trained and fine-tuned on A800-80G GPU clusters555We also utilize clusters of other types of GPUs, for example, Nvidia V100-32G, L20, L40 and etc. using DeepSpeed’s Stage-2 strategy with mixed precision (FP16/FP32) or BF16 (Rasley et al., 2020). We employ the AdamW optimizer (Loshchilov & Hutter, 2019) with a learning rate scheduler. To maximize computational efficiency in pre-training stage, we pack multiple graph sequences into single entries, optimizing context window utilization (Raffel et al., 2020) for certain graph datasets. Dataset-specific configurations are detailed in their respective sections below.
The variance is inherently low for most large-scale datasets (e.g., PCQM4Mv2, ogbl-ppa), where reporting variance is standard practice only when significant. For these datasets, 3–5 runs consistently yield minimal variance (as shown in tables). For the Triangles dataset, variance is higher—particularly on out-of-distribution (OOD) test data. So we conducted 10 runs to ensure robustness.
The pre-training and fine-tuning paradigm is illustrated in Fig. 2.